OpenStudy (anonymous):

6.Describe how to translate the graph of y=sqrt{x} to obtain the graph of y=sqrt{x-4}. shift down 4 units shift left 4 units shift up 4 units shift right 4 units

OpenStudy (anonymous):

@melodycrazysouthern

OpenStudy (anonymous):

21.Describe how to translate the graph of y=sqrt{x} to obtain the graph of y=sqrt{x}+5. shift up 5 unit shift left 5 unit shift right 5 unit shift down 5 unit

OpenStudy (anonymous):

@ganeshie8 6 and 21

ganeshie8 (ganeshie8):

changing f(x) to f(x-k) shifts the graph to "right" by "k" units

ganeshie8 (ganeshie8):

based on that, can you guess the answer for #6 ?

OpenStudy (anonymous):

C?

ganeshie8 (ganeshie8):

nope

OpenStudy (anonymous):

B?

ganeshie8 (ganeshie8):

nope, are you using the hint i provided you ?

OpenStudy (anonymous):

looked at the wrong problem. D

ganeshie8 (ganeshie8):

Yes! y = sqrt(x-4) shifts the graph of y = sqrt(x) to right by 4 units

OpenStudy (anonymous):

ok. thank you. Can you help me with 21

ganeshie8 (ganeshie8):

here i have graphed y = sqrt(x) https://www.desmos.com/calculator/ixniqvz9rr

ganeshie8 (ganeshie8):

click that link and enter y = sqrt(x) + 5 on the left side you will get another graph

ganeshie8 (ganeshie8):

compare them

ganeshie8 (ganeshie8):

Here they are :